Moonscape has just completed an initial public offering. The firm sold 4 million shares at an offer price of $12 per share. The underwriting spread was $0.40 a share. The price of the stock closed at $18.00 per share at the end of the first day of trading. The firm Incurred $400,000 in legal, administrative, and other costs. What were flotation costs as a fraction of funds raised? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.) 


Costs as percent of funds raised = _______ %
Answer #1

Funds raised = Shares*offer price

= 4000000*12

= 48,000,000

Costs = Shares*underwriting spread +Shares*underpricing amount + legal, administrative and other costs

= 4000000*0.4+4000000*(18-12)+ 400000

= 26,000,000

Costs as a percentage of funds raised = 26000000/48000000

= 54.17%